OpenStudy (anonymous):

how do you solve x - 30% = 22.40?? I know the answer is 32.00 but how did i arrive here???

OpenStudy (lgbasallote):

i think you mean x - 30%x = 22.40

OpenStudy (lgbasallote):

first thing you need to do is change 30% into a decimal \[\implies x - 0.30x = 22.40\] got it?

OpenStudy (anonymous):

yes, and...

OpenStudy (lgbasallote):

then remember your algebra... x - 0.30 x is like (1 - 0.30)x 1 - 0.30 is 0.70 so x - 0.30x = 0.70x \[\implies 0.70x = 22.40\] still following?

OpenStudy (anonymous):

yes, ty

OpenStudy (lgbasallote):

then the final step would be to divide both sides by 0.70 \[\implies x = \frac{22.40}{0.70}\] this gives you 32.00 like you said

OpenStudy (anonymous):

thank you. G

OpenStudy (lgbasallote):

welcome